Suppose that the regression for predicting weight (in pounds) from Height (in Select one answer inches) is given by Weight =-115 + 3.6(Height) Which of the following statements is correct? I. A person who is 61 inches tall will weigh 104.6 pounds II. For every additional inch of height, the predicted weight will increase, on average, by 3.6 pounds. III. The correlation between weight and height is negative. ı points A. I only B. II only C. III only D. O I and II only E. II and III only
Answer #1

Weight corresponding to 61 inches = -115 + 61x3.6 = 104.6 pounds

Therefore, statement I is correct

Weight = -115 + 3.6 x Height

Statement II is correct

The correlation between weight and height is positive because with increase in height, weight also increases. Statement II is incorrect.

Ans: D. I and II only